OpenStudy (anonymous):

(2-x^2)y'' + xy' + 16y = 0 and y'' - x^2y' - 3xy = 0 Find the general form in powers of x, stating the recurrence relation and radius of convergence.

OpenStudy (amistre64):

should i assume this relates to using power series?

OpenStudy (anonymous):

yes it is

OpenStudy (amistre64):

and is it a system of equations that must be satisfied as opposed to simply 2 separate equations?

OpenStudy (anonymous):

Sorry I posted 2 separate questions in 1 that I should have posted as 2.

OpenStudy (amistre64):

whew!! so it is spose to be 2 separate equations .... thats a relief

OpenStudy (amistre64):

granted it was more interesting as a system, but at leasts its simpler now :)

OpenStudy (amistre64):

assume theres a power series that works; y = :(0,inf): an x^n y' = :(1,inf): an n x^(n-1) y'' = :(2,inf): an n(n-1) x^(n-2) and plug them into the eq

OpenStudy (amistre64):

y'' - x^2y' - 3xy = 0 y'' = :(2,inf): an n(n-1) x^(n-2) -x^2 y' = :(1,inf): -an n x^(n+1) -3x y = :(0,inf): -3an x^(n+1) we want them all to start at the same summation and exponents for good measure

OpenStudy (amistre64):

:(2,inf): an n(n-1) x^(n-2) :(1+3,inf): -a(n-3) (n-3) x^(n+1-3) :(0+3,inf): -3a(n-3) x^(n+1-3) :(2-2,inf): an n(n-1+2) x^(n-2+2) :(4-2,inf): -a(n-3+2) (n+2-3) x^(n-2+2) :(3-2,inf): -3a(n-3+2) x^(n-2+2) :(0,inf): a(n+2) (n+2)(n+1) x^(n) :(2,inf): -a(n-1) (n-1) x^(n) :(1,inf): -3a(n-1) x^(n) you think thats good so far?

OpenStudy (amistre64):

2 a2 + 6x a3 + :(2,inf): a(n+2) (n+2)(n+1) x^(n) :(2,inf): -a(n-1) (n-1) x^(n) -3x a0 + :(2,inf): -3a(n-1) x^(n) with any luck im remembering this correctly

OpenStudy (amistre64):

a0,a2,a3 = 0 [a(n+2) (n+2)(n+1) -a(n-1) (n-1) -3a(n-1)] = 0 i spose i shoulda left an an in there someplace, dontcha think?

OpenStudy (amistre64):

im going to paper for this

OpenStudy (amistre64):

http://archive.org/details/CalculusRevisitedPart3-PowerSeriesSolutions while im dillydallying about, this is a good review

OpenStudy (amistre64):

:(2,inf): an n(n-1) x^(n-2) :(1+3,inf): -a(n-3) (n-3) x^(n+1-3) :(0+3,inf): -3a(n-3) x^(n+1-3) :(2,inf): an n(n-1) x^(n-2) :(4,inf): -a(n-3) (n-3) x^(n-2) :(3,inf): -3a(n-3) x^(n-2) 2 a2 + 6 a3 x + :(4,inf): an n(n-1) x^(n-2) :(4,inf): -a(n-3) (n-3) x^(n-2) - 3 a0 x + :(4,inf): -3a(n-3) x^(n-2) 2 a2 = 0 when a2 = 0 6 a3 - 3 a0 = 0 :(4,inf): [ an n(n-1) - a(n-3) (n-3) -3a(n-3) ] x^(n-2) = 0 an n(n-1) = a(n-3) [ (n-3) + 3] an = a(n-3) /(n-1) we have a recurrsive equation such that : \[a_2=0;\ a_0=2a_3\] \[a_n=\frac{a_{n-3}}{n-1}\] \[a_3=\frac{a_0}{2};\ a_6=\frac{a_0}{2*5};\ a_9=\frac{a_0}{2*5*8}...\] \[a_4=\frac{a_1}{3};\ a_7=\frac{a_1}{3*6};\ a_{10}=\frac{a_1}{3*6*9}...\] \[a_5=\frac{a_2}{4};\ a_8=\frac{a_2}{4*7};\ a_{11}=\frac{a_2}{4*7*11}...=0\] which is all i think i can do for that; not real sure how to wrap it up tho